# Leading Higher Derivative Corrections to Multipole Moments of Kerr-Newman Black Hole

### v2

@@ 1 Introduction
Similar to black hole thermodynamics, black hole multipole moments will also receive higher-derivative corrections, except for the mass monopole and the current dipole moment, corresponding to total mass and angular momentum that are typically chosen to be fixed under the perturbation. Black hole multipole moments play a crucial role in describing the external fields and gravitational wave radiation, potentially offering a new window into the footprints of quantum gravity [5, 6, 7, 8] beyond standard General Relativity. Upcoming observations of gravitational waves from extreme mass-ratio inspirals (EMRIs) [9, 10, 11, 12, 13, 14] may reveal whether black hole mass and spin multipole moments match predictions from classical general relativity or suggest modifications induced by new physics beyond the standard model. It is thus urgent to understand in an effective theory of gravity, whether black hole multipole moments are meaningful physical observables, i.e. invariant under field redefinitions. Based on traditional approach, to obtain black hole multipole moments, one needs to first solve for higher derivative corrections to rotating black holes which is technically quite difficult. Until now, there is no proof that in a generic theory of gravity with higher derivative corrections, the black hole multipole moments are invariant under field redefinitions, except for very few specific examples. For instance, the recent work [15] has computed the multipole moments of Kerr black holes in pure gravity with cubic curvature corrections and showed that they are invariant under field redefinitions of the metric utilizing the Ricci flatness of Kerr solution. For non Ricci flat solutions, such as Kerr-Newman black holes, the situation is unknown. In fact, it was speculated in [15] that for non-Ricci-flat solutions, higher derivative corrections to multipole moments might be affected by field redefinitions.
@@
At the first order in 4-derivative couplings, corrections from the parity even and odd terms to the black hole solution decouple from each other. Thus we can analyze these two cases separately. When parity preserving 4-derivative terms are switched on, we are able to obtain the perturbed solution up to ${\cal O}(\chi^{7})$, while concerning only parity odd 4-derivative interactions, we can reach ${\cal O}(\chi^{8})$. From the approximate solution, we could read off corrections to both the gravitational and electromagnetic multipole moments at first few levels. Similar to the pure gravity case, parity preserving 4-derivative interactions only modify multipole moments $\{M_{2n},\mathcal{S}_{2n+1},\mathcal{Q}_{2n},\mathcal{P}_{2n+1}\}$ that are already nonzero at the leading order. In the parity odd case, the 4-derivative interactions contribute to multipole moments $\{M_{2n+1},\mathcal{S}_{2n},\mathcal{Q}_{2n+1},\mathcal{P}_{2n}\}$ that vanish at the leading order. Thus their very presence breaks the equatorial symmetry of the solution and has significant observational implications. Most importantly, we show that these results can be expressed in a way that is manifestly invariant under field redefinitions.
@@ 2 Multipole moments of Kerr-Newman black hole in 2-derivative theory
$g_{\varphi\varphi}$ $=$ $r^{2}(1-x^{2})\left[1+\sum_{\ell\geq 0}^{\infty}\frac{1}{r^{\ell+1}}\sum_{\ell^{\prime}\leq\ell}c_{\ell\ell^{\prime}}^{(\varphi\varphi)}P_{\ell^{\prime}}\right],\quad g_{rx}=r\left[\sum_{\ell\geq 0}^{\infty}\frac{1}{r^{\ell+1}}\sum_{\ell^{\prime}\leq\ell}c_{\ell\ell^{\prime}}^{(rx)}P_{\ell^{\prime}}\right]\ ,$ (7)
@@ 3.1 Parity-even case
$r^{(1)}(r_{S},x_{S})=\sum_{k=-1}^{\infty}\sum_{p=0}^{k+1}b_{k,p}\frac{x_{S}^{p}}{r_{S}^{k}}\ ,\qquad x^{(1)}(r_{S},x_{S})=(1-x_{S}^{2})\sum_{k=0}^{\infty}\sum_{p=0}^{k+1}c_{k,p}\frac{x_{S}^{p}}{r_{S}^{k}}\ .$ (34)
The coefficients $b_{k,p}$ and $c_{k,p}$ are fixed by requiring the modified metric to remain in the standard form (7), from which we read off corrections to the gravitational multipole moments due to the parity even 4-derivative interactions
$\delta M_{0}^{(e)}$ $=$ $\frac{1}{35\mu}\Big{[}\big{(}\alpha_{0}-24\alpha_{1}-4\alpha_{3}\big{)}\chi_{Q}^{2}-\alpha_{2}(35\chi_{a}^{2}+4\chi_{Q}^{2})\Big{]}+\frac{1}{13860\mu}\Big{[}1386\alpha_{2}\chi_{a}^{4}$
$+3(65\alpha_{0}-1252\alpha_{1}-3201\alpha_{2}-337\alpha_{3})\chi_{a}^{2}\chi_{Q}^{2}-44(3\alpha_{0}+4\alpha_{2}+24\alpha_{1}+4\alpha_{3})\chi_{Q}^{4}\Big{]}$
$+\frac{1}{720720\mu}\Big{[}27027\alpha_{2}\chi_{a}^{6}-6(2148\alpha_{0}-8652\alpha_{1}+883\alpha_{2}-1299\alpha_{3})\chi_{a}^{4}\chi_{Q}^{2}$
$+(5999\alpha_{0}-149128\alpha_{1}-339144\alpha_{2}-46562\alpha_{3})\chi_{a}^{2}\chi_{Q}^{4}-1040(9\alpha_{0}-4\alpha_{2}-24\alpha_{1}$
$-4\alpha_{3})\chi_{Q}^{6}\Big{]}+\frac{2(c_{2}+16c_{7})\chi_{Q}^{2}}{45045\mu}\Big{[}290\chi_{a}^{2}\chi_{Q}^{2}+162\chi_{a}^{4}+65\chi_{Q}^{4}+117\chi_{a}^{2}-143\chi_{Q}^{2}$
[…]
$\delta M^{(e)}_{2}$ $=$ $2\alpha_{2}\mu\chi_{a}^{2}+\frac{\mu\chi_{a}^{2}}{210}\Big{[}168\alpha_{2}\chi_{a}^{2}-(13\alpha_{0}-284\alpha_{1}-325\alpha_{2}-59\alpha_{3})\chi_{Q}^{2}\Big{]}$
$-\frac{\mu\chi_{a}^{2}}{138600}\Big{[}24255\alpha_{2}\chi_{a}^{4}-30(67\alpha_{0}+1384\alpha_{1}+3630\alpha_{2}+370\alpha_{3})\chi_{a}^{2}\chi_{Q}^{2}+22(213\alpha_{0}$
$-4176\alpha_{1}-5393\alpha_{2}-1004\alpha_{3})\chi_{Q}^{4}\Big{]}-\frac{2(c_{2}+16c_{7})\mu\chi_{a}^{2}\chi_{Q}^{2}}{3465}(9\chi_{a}^{2}-11\chi_{Q}^{2}-99)\ ,$
$\delta M^{(e)}_{4}$ $=$ $-4\alpha_{2}\mu^{3}\chi_{a}^{4}-\frac{\mu^{3}\chi_{a}^{4}}{3675}\Big{[}\alpha_{2}(2205\chi_{a}^{2}+10811\chi_{Q}^{2})-35(10\alpha_{0}-212\alpha_{1}-47\alpha_{3})\chi_{Q}^{2}\Big{]}$
$-\frac{2\mu^{3}\chi_{a}^{4}\chi_{Q}^{2}}{35}(c_{2}+16c_{7})\ ,$
$\delta M^{(e)}_{6}$ $=$ $6\alpha_{2}\mu^{5}\chi_{a}^{6}\ ,$
$\delta\mathcal{S}_{1}^{(e)}$ $=$ $-\alpha_{2}\chi_{a}-\frac{\chi_{a}}{420}\Big{[}\alpha_{2}(378\chi_{a}^{2}+139\chi_{Q}^{2})-(19\alpha_{0}-428\alpha_{1}-83\alpha_{3})\chi_{Q}^{2}\Big{]}$
$+\frac{\chi_{a}}{55440}\Big{[}7623\alpha_{2}\chi_{a}^{4}-6(2\alpha_{0}+2636\alpha_{1}+7293\alpha_{2}+707\alpha_{3})\chi_{a}^{2}\chi_{Q}^{2}-22(3\alpha_{0}+612\alpha_{1}$
$+242\alpha_{2}+137\alpha_{3})\chi_{Q}^{4}\Big{]}+\frac{\chi_{a}}{5765760}\Big{[}324324\alpha_{2}\chi_{a}^{6}-6(24477\alpha_{0}+31517\alpha_{2}-46908\alpha_{1}$
$-4815\alpha_{3})\chi_{a}^{4}\chi_{Q}^{2}-(2201\alpha_{0}+1594568\alpha_{1}+3531684\alpha_{2}+472882\alpha_{3})\chi_{a}^{2}\chi_{Q}^{4}-260(189\alpha_{0}$
$+1212\alpha_{1}+763\alpha_{2}+367\alpha_{3})\chi_{Q}^{6}\Big{]}+\frac{2(c_{2}+16c_{7})\chi_{a}\chi_{Q}^{2}}{45045}\Big{[}290\chi_{a}^{2}\chi_{Q}^{2}+162\chi_{a}^{4}+65\chi_{Q}^{4}$
$+117\chi_{a}^{2}-143\chi_{Q}^{2}-1287\Big{]}\ ,$
$\delta\mathcal{S}_{3}^{(e)}$ $=$ $3\alpha_{2}\mu^{2}\chi_{a}^{3}+\frac{\mu^{2}\chi_{a}^{3}}{700}\Big{[}490\alpha_{2}\chi_{a}^{2}-(55\alpha_{0}-1180\alpha_{1}-1179\alpha_{2}-255\alpha_{3})\chi_{Q}^{2}\Big{]}$
$-\frac{\mu^{2}\chi_{a}^{3}}{1940400}\Big{[}412335\alpha_{2}\chi_{a}^{4}-42(39369\alpha_{2}+1330\alpha_{0}+14500\alpha_{1}+3865\alpha_{3})\chi_{a}^{2}\chi_{Q}^{2}+22(3201\alpha_{0}$
$-71916\alpha_{1}-86188\alpha_{2}-17419\alpha_{3})\chi_{Q}^{4}\Big{]}+\frac{2(c_{2}+16c_{7})\mu^{2}\chi_{a}^{3}\chi_{Q}^{2}}{3465}(11\chi_{Q}^{2}-9\chi_{a}^{2}+99)\ ,$
$\delta\mathcal{S}_{5}^{(e)}$ $=$ $-5\alpha_{2}\mu^{4}\chi_{a}^{5}-\frac{\mu^{4}\chi_{a}^{5}}{2940}\Big{[}\alpha_{2}(1470\chi_{a}^{2}+8957\chi_{Q}^{2})-7(47\alpha_{0}-988\alpha_{1}-223\alpha_{3})\chi_{Q}^{2}\Big{]}$
$-\frac{2\mu^{4}\chi_{a}^{5}\chi_{Q}^{2}}{35}(c_{2}+16c_{7})\ ,$
$\delta\mathcal{S}_{7}^{(e)}$ $=$ $7\alpha_{2}\mu^{6}\chi_{a}^{7}\ .$ (51)
Thus it is evident that the parity even 4-derivative terms will not contribute nontrivially to $M_{2n+1}$ and ${\cal S}_{2n}$. In the current choice of integration constants, we noticed that the total mass and angular momentum seem to receive corrections and some of the multipole moments depend on 4-derivative couplings that are variant under field redefinitions. These terms are all proportional to $c_{2}+16c_{7}$.
In fact, the dependence of gravitational multipole moments on $c_{2}+16c_{7}$ is an arifact due to our bad choice of integration constants. The appropriate choice should be that the mass, angular momentum and the electric charge should receive no 4-derivative correction. In terms of new integration constants $\mu^{\prime},\,\chi_{a}^{\prime},\,\chi_{Q}^{\prime}$ defined below
@@
where $\delta\mu,\,\delta\chi_{a},\,\delta\chi_{Q}$ are given in Appendix B, we see that the mass, angular momentum, and electric charge of the Kerr-Newman black hole become independent of 4-derivative couplings. Meanwhile, all the terms that are variant under field redefinitions disappear from the corrections to multipole moments. Below we present the multipole moments written in terms of the new integration constants, and for the tidiness, we remove the “prime” from notation. Using the approximate solution, we are able to obtain the first few gravitational multipole moments listed below
@@
$\delta\mathcal{S}_{3}^{(e)}$ $=$ $\frac{23}{25}\alpha_{2}\mu^{2}\chi_{a}^{3}\chi_{Q}^{2}-\frac{\mu^{2}\chi_{a}^{3}\chi_{Q}^{2}}{4900}\Big{[}\alpha_{2}(637\chi_{a}^{2}-3501\chi_{Q}^{2})+(102\alpha_{0}-1172\alpha_{1}-293\alpha_{3})\chi_{Q}^{2}\Big{]}\ ,$ (53)
@@
$\delta\mathcal{Q}_{6}^{(e)}$ $=$ $\frac{125\alpha_{2}\mu^{5}\chi_{a}^{6}\chi_{Q}}{1386}\ ,$ (55)
@@
The results depends only on the invariant combinations $(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})$ of coupling constants defined in (22).
@@ 3.2 Parity-odd case
$\mathcal{L}_{4}^{(o)}=d_{1}R_{\mu\nu}F^{\mu\rho}\widetilde{F}^{\nu}_{\ \rho}+d_{2}R_{\mu\nu\rho\sigma}F^{\mu\nu}\widetilde{F}^{\rho\sigma}+d_{3}\widetilde{F}_{\mu\nu}F^{\mu\nu}F^{2}+d_{4}\widetilde{F}_{\mu\nu}F^{\nu\rho}F_{\rho\sigma}F^{\sigma\mu}\ ,$ (60)
where $\widetilde{F}_{\mu\nu}:=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ and we have used the useful identity that $g_{\mu\nu}F_{\rho\sigma}\widetilde{F}^{\rho\sigma}=4F_{\mu\rho}\widetilde{F}_{\nu}^{\ \rho}$. Because of the same identity, the parity odd field redefinition of the metric has only one structure
$g_{\mu\nu}\rightarrow g_{\mu\nu}+\lambda_{5}F_{\mu\rho}\widetilde{F}_{\nu}{}^{\rho}\,,$ (61)
which shifts the coupling constants in (60) according to as
$d_{1}\rightarrow d_{1}+\lambda_{5},\quad d_{2}\rightarrow d_{2},\quad d_{3}\rightarrow d_{3}-\frac{\lambda_{5}}{8},\quad d_{4}\rightarrow d_{4}+\frac{\lambda_{5}}{2}\ .$ (62)
One can easily check that the combinations of 4-derivative couplings below are invariant under (62)
$\beta_{0}=4d_{3}+d_{4},\quad\beta_{1}=d_{2},\quad\beta_{2}=8d_{3}+d_{1}\ .$ (63)
@@
$\delta\mathcal{Q}_{7}^{(o)}$ $=$ $\frac{121388\beta_{1}\mu^{6}\chi_{a}^{7}\chi_{Q}}{225225}\ ,$ (67)
@@
It is evident that all the expressions above are manifestly invariant under field redefinitions. Note that the magnetic charge ${\cal P}_{0}$, which is zero in our original Kerr-Newman black hole, remains uncorrected, whilst the magnetic higher even multipole moments all receive corrections by the party-odd 4-derivative terms.
@@ 4 Conclusion
Black hole multipole moments are potentially useful observables for probing the nonlinear structures of General Relativity and its modifications due to unknown UV physics. In this work, we made progress in addressing a key question: do higher derivative perturbative corrections to black hole multipole moments are invariant under field redefinitions that are not supposed to affect physical quantities in a low energy effective theory of quantum gravity. So far, traditional methods for calculating multipole moments have not explicitly demonstrated this property for general effective theory of gravity. As far as we are aware, the field redefinition invariance of black hole multipole moments has only been shown for $D=4$ Einstein gravity extended by cubic curvature terms [15], where the Ricci flatness of the leading order solution had played a role in the proof. Here, we further investigated the leading higher-derivative corrections (19,60) to the non-Ricci flat Kerr-Newman black hole (4). We find that the leading higher derivative corrections to black hole multipole moments indeed depend only on the combinations of coupling constants that are inert under field redefinitions.
@@
In $D=4$, the leading higher derivative corrections to Einstein-Maxwell theory are classified by their properties under parity transformations. For parity preserving 4-derivative interactions, we are able to obtain the perturbed solution up to ${\cal O}(\chi^{7})$. Similar to the pure gravity case, 4-derivative interactions only add corrections to multipole moments that are actually nonzero at the leading order, namely, $\{M_{2n+2},\mathcal{S}_{2n+3},\mathcal{Q}_{2n+2},\mathcal{P}_{2n+1}\}$ for $n\geq 0$. Concerning the parity odd 4-derivative interactions, we obtain the perturbed solution up to ${\cal O}(\chi^{8})$. We find that while these higher-derivative terms do not contribute to black hole thermodynamics, they affect the perturbative solutions, thereby modifying the black hole multipole moments. Notably, they contribute to multipole moments $\{M_{2n+1},\mathcal{S}_{2n},\mathcal{Q}_{2n-1},\mathcal{P}_{2n}\}$ for $n\geq 1$ that all vanish at the leading order. This feature has significant observational implications. If they are observed, it would indicate the presence of parity odd higher-derivative corrections to General Relativity.

### v3

@@ 1 Introduction
Similar to black hole thermodynamics, black hole multipole moments will also receive higher-derivative corrections, except for the mass monopole and the current dipole moment, corresponding to total mass and angular momentum that are typically chosen to be fixed under the perturbation. Black hole multipole moments play a crucial role in describing the external fields and gravitational wave radiation, potentially offering a new window into the footprints of quantum gravity [5, 6, 7, 8] beyond standard General Relativity. Upcoming observations of gravitational waves from extreme mass-ratio inspirals (EMRIs) [9, 10, 11, 12, 13, 14] may reveal whether black hole mass and spin multipole moments match predictions from classical general relativity or suggest modifications induced by new physics beyond the standard model. It is thus urgent to understand in an effective theory of gravity, whether black hole multipole moments are meaningful physical observables, i.e. invariant under field redefinitions. Based on traditional approach, to obtain black hole multipole moments, one needs to first solve for higher derivative corrections to rotating black holes which is technically quite difficult. Until now, there is no proof that in a generic theory of gravity with higher derivative corrections, the black hole multipole moments are invariant under field redefinitions, except for very few specific examples. For instance, the recent work [15] has computed the multipole moments of Kerr black holes in pure gravity with cubic curvature corrections and showed that they are invariant under field redefinitions of the metric utilizing the Ricci flatness of Kerr solution. For non Ricci flat solutions, such as Kerr-Newman black holes, the situation is unknown.
@@
At the first order in 4-derivative couplings, corrections from the parity even and odd terms to the black hole solution decouple from each other. Thus we can analyze these two cases separately. When parity preserving 4-derivative terms are switched on, we are able to obtain the perturbed solution up to ${\cal O}(\chi^{7})$, while concerning only parity odd 4-derivative interactions, we can reach ${\cal O}(\chi^{8})$  [^1] Here we can only reach ${\cal O}(\chi^{7})$ or ${\cal O}(\chi^{8})$, which seems low compared to the vacuum Kerr case [17]. This is due to the limited computation power and the further complexity of the field equations caused by more variables and higher derivative terms switched on. . From the approximate solution, we could read off corrections to both the gravitational and electromagnetic multipole moments at first few levels. Similar to the pure gravity case, parity preserving 4-derivative interactions only modify multipole moments $\{M_{2n},\mathcal{S}_{2n+1},\mathcal{Q}_{2n},\mathcal{P}_{2n+1}\}$ that are already nonzero at the leading order. In the parity odd case, the 4-derivative interactions contribute to multipole moments $\{M_{2n+1},\mathcal{S}_{2n},\mathcal{Q}_{2n+1},\mathcal{P}_{2n}\}$ that vanish at the leading order. Thus, their very presence breaks the equatorial symmetry of the solution and has significant observational implications (see [18] for a recent study and references therein). Most importantly, we show that these results can be expressed in a way that is manifestly invariant under field redefinitions.
@@ 2 Multipole moments of Kerr-Newman black hole in 2-derivative theory
$g_{\varphi\varphi}$ $=$ $r^{2}(1-x^{2})\left[1+\sum_{\ell\geq 0}^{\infty}\frac{1}{r^{\ell+1}}\sum_{\ell^{\prime}\leq\ell}c_{\ell\ell^{\prime}}^{(\varphi\varphi)}P_{\ell^{\prime}}\right],\quad g_{rx}=r\left[\sum_{\ell\geq 0}^{\infty}\frac{1}{r^{\ell+1}}\sum_{\ell^{\prime}\leq\ell}c_{\ell\ell^{\prime}}^{(rx)}P_{\ell^{\prime}}^{\prime}\right]\ ,$ (7)
@@ 3.1 Parity-even case
It is interesting to note that the following structures are independently invariant under the electromagnetic duality transformation [25]
$F^{\mu}{}_{\nu}F^{\nu}{}_{\rho}F^{\rho}{}_{\sigma}F^{\sigma}{}_{\mu}-\frac{1}{4}(F^{2})^{2},\quad R^{\mu\nu}F_{\mu\rho}F_{\nu}{}^{\rho}-\frac{1}{4}RF^{2}\ ,$
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2},\quad R_{\mu\nu}R^{\mu\nu},\quad R^{2}\ .$ (20)
@@
$r^{(1)}(r_{S},x_{S})=\sum_{k=-1}^{\infty}\sum_{p=0}^{k+1}b_{k,p}\frac{x_{S}^{p}}{r_{S}^{k}}\ ,\qquad x^{(1)}(r_{S},x_{S})=(1-x_{S}^{2})\sum_{k=1}^{\infty}\sum_{p=0}^{k-1}c_{k,p}\frac{x_{S}^{p}}{r_{S}^{k}}\ .$ (35)
The coefficients $b_{k,p}$ and $c_{k,p}$ are fixed by requiring the modified metric to remain in the standard form (7), from which we read off corrections to the gravitational multipole moments due to the parity even 4-derivative interactions. In terms of new integration constants $\mu^{\prime},\,\chi_{a}^{\prime},\,\chi_{Q}^{\prime}$ defined below
@@
where $\delta\mu,\,\delta\chi_{a},\,\delta\chi_{Q}$ are given in Appendix B, we find that the mass, angular momentum, and electric charge of the Kerr-Newman black hole become independent of 4-derivative couplings. Meanwhile, all the multipole moments are manifestly invariant under field redefinitions. We also note that corrections to mass multipole moments behave as $\delta M_{n}\sim\chi^{n+2}$ (+higher order in $\chi$). Therefore, in the parity-even case, with the perturbed solution up to ${\cal O}(\chi^{7})$, we can only obtain non-trivial corrections up to $M_{4}$.
Below we present the 4-derivative corrections to multipole moments written in terms of the new integration constants, and for the tidiness, we remove the “prime” from notation. Using the approximate solution, we are able to obtain the first few gravitational multipole moments listed below
@@
$\delta\mathcal{S}_{3}^{(e)}$ $=$ $\frac{23}{25}\alpha_{2}\mu^{2}\chi_{a}^{3}\chi_{Q}^{2}-\frac{\mu^{2}\chi_{a}^{3}\chi_{Q}^{2}}{4900}\Big{[}\alpha_{2}(637\chi_{a}^{2}-3501\chi_{Q}^{2})+(102\alpha_{0}-1172\alpha_{1}-293\alpha_{3})\chi_{Q}^{2}\Big{]}\ ,$
@@
$\delta\mathcal{Q}_{6}^{(e)}$ $=$ $\frac{125\alpha_{2}\mu^{5}\chi_{a}^{6}\chi_{Q}}{1386}\ ,$
@@
The results depends only on the invariant combinations $(\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3})$ of coupling constants defined in (23). It is evident that the parity even 4-derivative terms will not contribute nontrivially to $M_{2n+1}$ and ${\cal S}_{2n}$.
@@ 3.2 Parity-odd case
$\mathcal{L}_{4}^{(o)}=d_{1}RF^{\mu\rho}\widetilde{F}_{\mu\rho}+d_{2}R_{\mu\nu\rho\sigma}F^{\mu\nu}\widetilde{F}^{\rho\sigma}+\frac{d_{3}}{4}\widetilde{F}_{\mu\nu}F^{\mu\nu}F^{2}\ ,$ (42)
where $\widetilde{F}_{\mu\nu}:=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}F^{\rho\sigma}$ and we have used the useful identity that $g_{\mu\nu}F_{\rho\sigma}\widetilde{F}^{\rho\sigma}=4F_{\mu\rho}\widetilde{F}_{\nu}^{\ \rho}$. Note that at 4-derivative level, there exists no parity-odd term preserving electromagnetic duality invariance. Because of the same identity, the parity odd field redefinition of the metric has only one structure
$g_{\mu\nu}\rightarrow g_{\mu\nu}+\lambda_{5}g_{\mu\nu}F_{\rho\sigma}\widetilde{F}^{\rho\sigma}\,,$ (43)
which shifts the coupling constants in (42) as
$d_{1}\rightarrow d_{1}+\lambda_{5},\quad d_{2}\rightarrow d_{2},\quad d_{3}\rightarrow d_{3}\ .$ (44)
It is evident that 4-derivative couplings below are invariant under the field redefinition
$\beta_{0}=d_{3},\quad\beta_{1}=d_{2}\ .$ (45)
@@
$\delta\mathcal{Q}_{7}^{(o)}$ $=$ $\frac{121388\beta_{1}\mu^{6}\chi_{a}^{7}\chi_{Q}}{225225}\ ,$
@@
It is evident that all the expressions above are manifestly invariant under field redefinitions. Note that the magnetic charge ${\cal P}_{0}$, which is zero in our original Kerr-Newman black hole, remains uncorrected, whilst the magnetic higher even multipole moments all receive corrections by the party-odd 4-derivative terms. Clearly, observation of the multipole moments above would indicate the presence of parity-odd terms and the breaking of electromagnetic duality invariance.
@@ 4 Conclusion
Black hole multipole moments are potentially useful observables for probing the nonlinear structures of General Relativity and its modifications due to unknown UV physics. In this work, we made progress in addressing a key question: are higher derivative perturbative corrections to black hole multipole moments invariant under field redefinitions. So far, traditional methods for calculating multipole moments have not explicitly demonstrated this property for general effective theory of gravity. As far as we are aware, the field redefinition invariance of black hole multipole moments has only been shown for $D=4$ Einstein gravity extended by cubic curvature terms [15], where the Ricci flatness of the leading order solution had played a role in the proof. Here, we further investigated the leading higher-derivative corrections (19,42) to the non-Ricci flat Kerr-Newman black hole (4). We find that the leading higher derivative corrections to black hole multipole moments indeed depend only on the combinations of coupling constants that are inert under field redefinitions.
@@
In $D=4$, the leading higher derivative corrections to Einstein-Maxwell theory are classified by their properties under parity transformations. For parity preserving 4-derivative interactions, we are able to obtain the perturbed solution up to ${\cal O}(\chi^{7})$. Similar to the pure gravity case, 4-derivative interactions only add corrections to multipole moments that are actually nonzero at the leading order, namely, $\{M_{2n+2},\mathcal{S}_{2n+3},\mathcal{Q}_{2n+2},\mathcal{P}_{2n+1}\}$ for $n\geq 0$. Concerning the parity odd 4-derivative interactions, we obtain the perturbed solution up to ${\cal O}(\chi^{8})$. We find that while these higher-derivative terms do not contribute to black hole thermodynamics, they affect the perturbative solutions, thereby modifying the black hole multipole moments. Notably, they contribute to multipole moments $\{M_{2n+1},\mathcal{S}_{2n},\mathcal{Q}_{2n-1},\mathcal{P}_{2n}\}$ for $n\geq 1$ that all vanish at the leading order. This feature has significant observational implications. If they are observed, it would indicate the presence of parity odd higher-derivative corrections to General Relativity [18].